OpenStudy (anonymous):

Alright. So you have a hemisphere with an 80 degree chunk taken out of it. The radius of the hemisphere is 6. You are given the mass 330 grams. You are trying to find the density.

OpenStudy (anonymous):

80/180=.444444444 80 degree chunk is 44% of the 1-.4444444=.55555556 so what is left of the hemishpere is 55.6% 330/.556=594g total mass of hemisphere is 594. volume of hemishpere is 4/3pi*6^3*1/2=452.4u^3 density is mass/volume=594g/452.4u^3 =1.313g/u^3